### v1

# SimBEV: A Synthetic Multi-Task Multi-Sensor Driving Data Generation Tool and Dataset

Abstract
Bird’s-eye view (BEV) perception for autonomous driving has garnered significant attention in recent years, in part because BEV representation facilitates the fusion of multi-sensor data. This enables a variety of perception tasks including BEV segmentation, a concise view of the environment that can be used to plan a vehicle’s trajectory. However, this representation is not fully supported by existing datasets, and creation of new datasets can be a time-consuming endeavor. To address this problem, in this paper we introduce SimBEV, an extensively configurable and scalable randomized synthetic data generation tool that incorporates information from multiple sources to capture accurate BEV ground truth data, supports a comprehensive array of sensors, and enables a variety of perception tasks including BEV segmentation and 3D object detection. We use SimBEV to create the SimBEV dataset, a large collection of annotated perception data from diverse driving scenarios.
1 Introduction
Autonomous driving promises a future with safer, cleaner, more efficient, and more reliable transportation systems [11, 58]. As the development of autonomous vehicle technology has accelerated in recent years, so has the need for perception algorithms capable of understanding complex driving scenarios in diverse environments [73, 7]. High-quality driving datasets have been at the center of the recent progress, serving as a foundation for training and benchmarking novel perception algorithms. It is essential for such datasets to encompass a large variety of driving scenarios and encapsulate a diverse set of road types, weather conditions, and traffic patterns, so perception models can effectively generalize to real-world situations [32, 53, 15].
As essential in this context is multimodal sensor fusion, which plays an important role in enhancing the performance of perception algorithms by compensating for weaknesses of one modality with the strengths of others [77, 5, 67]. Sensor fusion improves an autonomous vehicle’s understanding of its surrounding environment [59] (especially in adverse weather conditions [1]), enables robust decision-making in dynamic scenarios [19, 50], and opens the door to perception models that can accomplish multiple tasks simultaneously [35, 43, 47, 20]. Consequently, it is imperative for driving datasets to support a wide array of sensors and various computer vision tasks (such as object detection, semantic segmentation, and tracking) to facilitate the development of multi-faceted perception systems that leverage the strength of multiple sensing modalities.
Bird’s-eye view (BEV) perception has garnered significant attention in recent years for two primary reasons [40]. First, BEV representation is conducive to the fusion of information from different modalities, perspectives, and agents, and extracted BEV features can be used for various perception tasks. Second, BEV segmentation offers a concise, geometrically accurate and semantically rich view of the surrounding environment that can be directly used by motion planning, behavior prediction, and control algorithms. These two factors have given rise to perception algorithms that utilize BEV representation for 3D object detection, BEV segmentation, or both [35, 20, 61, 66, 30, 21, 23, 27, 72, 65, 41, 57, 70, 3, 33, 42, 14, 36, 17, 60, 28, 39, 26, 75, 10, 13, 45, 64, 69, 25, 74].
Figure 1: A data sample generated by SimBEV. The left half depicts a 360-degree view of the ego vehicle’s surroundings using different camera types (from top to bottom RGB, sematic segmentation, instance segmentation, depth, and optical flow cameras, respectively). On the right half, views of lidar, semantic lidar, radar, and the BEV ground truth are shown from top to bottom, respectively. Some images also contain 3D object bounding boxes.
However, despite the growing interest in BEV perception, few existing datasets support BEV segmentation. In the ones that do, either the BEV ground truth is limited to static map elements (drivable area, pedestrian crossing, etc.) [2], or BEV ground truth is only provided for a small window around the ego vehicle, obtained through combining map elements with select 3D object bounding boxes [16].
Creating a new dataset to address this gap is a challenging, time-consuming, and resource-intensive endeavor. Real-world driving data has to be collected over a long period of time to ensure weather conditions and traffic patterns present in the dataset are diverse, and requires (at least partly) labor-intensive hand annotation [56, 32]. As for synthetic datasets, they often consist of carefully designed scenarios that may not capture the full diversity of the driving environment, and may not be as challenging for perception algorithms as their creators think they are. Aside from all this, simply getting an overhead view of the ego vehicle in either case may not result in an accurate BEV ground truth due to the presence of vegetation and other structures (traffic lights, bridges, etc.) that obstruct the overhead view [31].
Dataset Year Scenes Annotated frames 2D Det 3D Det 2D Seg 3D Seg BEV Seg
Real-world KITTI [12] 2012 22 41K - - -
Cityscapes [6] 2016 - 25K - -
Mapillary [44] 2017 - 25K - - - -
ApolloScape [18] 2018 103 144K -
Argoverse [4] 2019 113 22K - - limited
Waymo Open [54] 2019 1150 230K - -
nuScenes [2] 2019 1000 40K limited
A*3D [46] 2020 - 39K - - - -
BDD100K [71] 2020 100K 12M - - -
Lyft Level 5 [16] 2021 366 46K - - - limited
Argoverse 2 [63] 2021 1000 6M - - limited
Synthetic SYNTHIA [49] 2016 - 13K - -
GTA-V [48] 2016 - 25K - - - -
ViPER [51] 2017 - 254K - -
All-in-One Drive [62] 2021 100 100K -
SHIFT [55] 2022 4850 2.5M - -
SimBEV 2024 320 102K
Table 1: Comparison of the size and supported tasks of the most notable existing single-vehicle driving datasets. SimBEV is the only dataset that provides full support for BEV perception.
To overcome these challenges, our paper makes two main contributions. First, we introduce SimBEV, an extensively configurable and scalable randomized synthetic data generation tool that incorporates information from multiple sources to capture accurate BEV ground truth data. In addition to diverse driving scenarios, SimBEV also provides 3D object bounding boxes and supports a variety of configurable perception and localization sensors. This makes it an invaluable tool for empowering researchers to explore various computer vision tasks, and ultimately, accelerate the development of more capable and resilient autonomous driving systems. Second, we use SimBEV to create the SimBEV dataset, a comprehensive large-scale dataset that can serve as a benchmark for a variety of perception tasks and have applications beyond it. One data sample generated by SimBEV is shown in Fig. 1.
2 Related Work
Real-world driving datasets generally target specific subsets of perception tasks, as the high costs associated with data collection and labeling limit their scope. One of the most prominent and oldest driving datasets is the KITTI dataset [12], which supports depth estimation and 2D and 3D object detection and tracking. Other notable image-based datasets include Cityscapes [6], and Mapillary [44], which are geared towards segmentation, while A*3D [46] focuses on 3D object detection. More recently, large-scale datasets such as BDD100K [71], Waymo Open [54], ApolloScape [18], Argoverse 2 [63], and nuScenes [2] have emerged, offering multi-modal data and multi-task annotations but primarily emphasizing object detection and tracking.
Synthetic driving datasets are compiled using graphics engines and physics simulators. For instance, SYNTHIA [49] incorporates RGB and semantically segmented images generated by its dedicated simulator. Video games have also served as a source of data. For example, GTA-V [48] offers RGB and semantically segmented images extracted from GTA. ViPER [51] expands upon it by including optical flow images and discrete environmental labels. The introduction of CARLA (CAR Learning to Act) Simulator [9], built upon Unreal Engine 4 (UE4), enabled the systemic generation of driving datasets. All-in-One Drive [62] is one such dataset, featuring support for multiple sensors and a focus on SPAD (Single-Photon Avalanche Detector)-lidars. Another notable dataset is the SHIFT dataset [55], a large-scale multi-task multi-modal synthetic dataset for autonomous driving, designed to simulate discrete and continuous shifts in weather and traffic conditions.
Existing datasets offer limited support for BEV segmentation. In nuScenes [2] BEV segmentation is only supported for static map elements (drivable area, lane line, pedestrian crossing, etc.). In Lyft Level 5 [16] and Argoverse 1 and 2 [4, 63], BEV ground truth is obtained through combining map elements and vehicle bounding boxes observable by the perception sensors, limiting BEV ground truth area and potentially missing some occluded objects.
Some vehicle-to-everything (V2X) datasets have support for BEV segmentation, but it is again limited. H-V2X [31] captures the BEV ground truth using overhead cameras installed along a 100 km highway, with the data limited to highway driving and mostly suitable for highway motion forecasting. The synthetic OPV2V [68] dataset provides BEV ground truth only for the drivable area, lane line, and vehicle classes, and the former can be inaccurate because it is calculated using waypoints only. Finally, the synthetic V2X-Sim [24] dataset captures the BEV ground truth using an overhead camera, limiting driving scenarios and causing inaccuracies due to vegetation, traffic light poles, and other structures that obstruct the overhead view.
3 SimBEV
SimBEV relies on a customized version of CARLA 0.9.15 [9] (see the Supplementary Material) to simulate the environment, perception sensors, and traffic behavior, although it is compatible with the standard release of CARLA as well. In the following, we describe how SimBEV operates and generates driving data.
3.1 Design
Instead of using carefully crafted scenarios, designing which can be very time-consuming, SimBEV relies on randomizing as many simulation parameters as possible to quickly create a statistically diverse set of scenarios. To create a new dataset, SimBEV generates and collects data from consecutive episodes, or scenes.
To start, the user configures the desired number of scenes for each map (i.e. driving environment, can be an existing CARLA map or a custom one) for each of the training, validation, and test sets. SimBEV checks to see if a SimBEV dataset already exists. If it does, SimBEV subtracts the number of existing scenes in that dataset from the number of desired scenes for each map. This feature allows the user to expand an already existing SimBEV dataset, or to continue with dataset creation in the event of a crash. SimBEV also allows the user to replace individual scenes.
At the start of each scene, SimBEV creates uniformly distributed waypoints within a certain distance of each other (specified by the user, 24.0 m by default) across the map’s roads and selects one at random to use to spawn the ego vehicle and the desired perception and localization sensors attached to it, though the user also has the option to specify a set of spawn points that SimBEV can choose from. SimBEV then configures the weather at random and, if at night, changes the intensity of street lights randomly.
To create background traffic, SimBEV randomly chooses the desired number of background vehicles and pedestrians, although the user has the option to specify each. It then uses the waypoints mentioned previously to spawn random vehicles at random positions, and spawns pedestrians at random on walkable areas of the map. CARLA’s traffic manager controls the behavior of vehicles and pedestrians throughout the episode.
Because all vehicles and pedestrians start from rest, SimBEV runs the simulation for a user-specified period of time (called warm-up duration, by default 4 s) before collecting data for a user-specified period of time (by default 16 s). During data collection, at each time step SimBEV collects and saves data from the desired sensors, and calculates and saves both 3D object bounding boxes and BEV ground truth masks. Following data collection, it saves meta-information about the data and a log of the scene, destroys the vehicles, pedestrians, and sensors, and starts a new scene.
3.2 Sensors
SimBEV supports a variety of sensors available in CARLA. They include five different camera types (RGB, semantic segmentation, instance segmentation, depth, and optical flow), two different lidar types (regular and semantic), radar, GNSS, and IMU, as shown in Fig. 1. The user has full control over each sensor’s parameters (e.g. a camera’s resolution or FoV), but the placement of the sensors is fixed for now. The cameras are placed at six locations above the vehicle similar to [2] to offer a 360-degree view of the vehicle’s surroundings, while a radar is placed on each side of the vehicle. GNSS and IMU sensors are placed at the center of the vehicle (the origin of the vehicle’s coordinate system), while the lidar is placed high above that center.
3.3 Scene configuration
As discussed in what follows, numerous parameters are randomized for each scene to ensure that scenes are as unique and diverse as possible.
Weather.
Weather in CARLA is configured using several parameters such as rain intensity, fog density, sun altitude angle, wind intensity, etc. By default, SimBEV randomly assigns these parameters for each scene to obtain diverse weather conditions, but the user has the option to set any of these parameters to a fixed value. For instance, setting the sun altitude angle to anything less than zero creates night-time scenes.
Traffic.
SimBEV randomly selects background vehicles from CARLA’s vehicle library, which includes sedans, vans, trucks, heavy goods vehicles (HGVs), buses, bicycles, motorcycles, and emergency vehicles (firetruck, ambulance, and police car). Emergency lights are randomly turned on for the latter vehicles, and the user can set the probability that this happens (by default 0.5). When possible, vehicle colors are randomly selected from a set of recommended colors, so e.g. a sedan can have one of 160 standard colors, but the firetruck will always be red. Additionally, some CARLA vehicles support articulated doors, so when these vehicles come to a stop (e.g. at a traffic light), SimBEV may randomly open one or all of their doors.
SimBEV randomly selects each vehicle’s maximum speed (relative to the speed limit, e.g. 10% over or under) and how close that vehicle can get to the vehicle in front of it when stopped. It also randomly selects how long each traffic light stays green. However, the user has the option to set any of these parameters to a fixed value.
SimBEV randomly selects pedestrians from CARLA’s walker library (which contains models of different ages, genders, races, and body types), sets their walking speed at random, and gives them random destinations to go to.
Lights.
SimBEV allows the user to turn off street and/or building lights (which are automatically turned on at night). Moreover, the user can choose to randomize building light colors, and/or change the intensity of all street lights by a fixed, if desired random, value. Moreover, based on the probability set by the user, SimBEV randomly turns off individual street lights to simulate broken street lights in the real world.
Reckless driving and jaywalking.
If desired by the user, some vehicles (including the ego vehicle) may drive recklessly, ignoring traffic lights, traffic signs, and collision with other vehicles and pedestrians, as shown in Fig. 2. The user can control the likelihood of reckless driving which can create interesting edge cases. Additionally, the user has control over the share of pedestrians who are allowed to (but may or may not) cross the road at any point, not just at crosswalks.
Figure 2: Reckless ego vehicle running over a cyclist.
3.4 Data annotations
SimBEV currently offers two main annotation types: 3D object bounding boxes and BEV ground truth. The output of some perception sensors (e.g. segmentation, depth, and optical flow cameras) can be used as annotation as well, but they are not discussed here.
3D object bounding boxes.
SimBEV currently collects 3D object bounding boxes that are within a user-configurable radius of the ego vehicle (by default 120.0 m) at each time step for the following six classes: car, truck (includes trucks, vans, HGVs, etc., but not buses), bus, motorcycle, bicycle, and pedestrian. SimBEV collects other object attributes alongside each bounding box such as the object’s ID, its linear and angular velocities, and its make, model, and color if the object is a vehicle. An optional post-processing step calculates the number of lidar and radar points that fall within each bounding box and adds a valid label to boxes with at least one point inside, invalid otherwise. This labeling is useful when training 3D object detection algorithms, as it can filter out objects that are not guaranteed to be visible by the perception sensors [2].
BEV segmentation.
Getting accurate BEV ground truth is the most critical aspect of SimBEV, which currently supports the following eight classes: road, car, truck, bus, motorcycle, bicycle, rider (human on a motorcycle or bicycle), and pedestrian. At each time step, the BEV ground truth is saved as a $C\times l\times l$ binary array, where $C$ is the number of classes and $l$ is the dimension of the BEV grid centered on the ego vehicle.
To calculate the BEV ground truth for non-road classes, we leverage the fact that ground elements in CARLA (roads, sidewalks, etc.) use one-way visible materials, i.e. they are see-through from the bottom, as shown in Fig. 3. This means that we can place a semantic segmentation camera 1 km above the ego vehicle facing down (to minimize perspective distortion) and another one 1 km below the ego vehicle facing up to catch what the overhead camera may miss because of obstructions. Both cameras have a $l\times l$ resolution and the FoV is calculated such that each pixel represents a $d\times d$ area on the ground. The ground truth for each non-road class is obtained by merging data from the two cameras using a logical or operation. By default, $l$ is set to 360 and $d$ is set to 0.4 m, creating a 144 m $\times$ 144 m box around the ego vehicle. This is larger than what is typically used for BEV segmentation (100 m $\times$ 100 m), but the additional area can help with data augmentation (rotation, translation, scaling) during training.
Figure 3: Ground elements in CARLA use one-way visible materials.
We follow an approach similar to [68] to get the ground truth for the road class. Specifically, we generate waypoints within a small distance of each other (specified by the user, we recommend setting it to $d$) across the map’s roads and note each waypoint’s lane width. We then calculate the mutual distance between these waypoints and the center of each cell of a $l\times l$ BEV grid centered on the ego vehicle where each cell represents a $d\times d$ area. For each grid cell, if a waypoint exists whose distance to the center of that cell is less than that waypoint’s lane width, that cell is labeled as road. Where our approach differs from [68] is that we then combine this information with data from the overhead camera and pass it through a binary closing operation to patch any potential holes, obtaining a more accurate ground truth.
Overall, this approach results in accurate BEV ground truth data and also lets us assign multiple labels to the same cell. For instance, a cell occupied by a rider on a motorcycle will have a rider label (obtained from the overhead camera), a motorcycle label (obtained from the underground camera), and a road label (obtained using the waypoints).
Our approach works everywhere except when roads with large elevation differences are present near the ego vehicle, e.g. when approaching an overpass. In those situations we do not use the semantic segmentation cameras. Instead, we only use the waypoint approach to calculate the road ground truth, and use 3D object bounding boxes to calculate the ground truth for other classes. While not as accurate as the previous method, the resulting ground truth is still acceptable. SimBEV intelligently switches between the two approaches when it detects two waypoints within 48.0 m of each other that have an elevation difference of at least 6.4 m, indicating they are on two different roads.
4 The SimBEV Dataset
To showcase SimBEV, we used it to generate the SimBEV dataset, a collection of 320 scenes spread across all 11 CARLA maps according to Tab. 2. Because Town12 and Town13 (CARLA’s two large maps) share many common features but use different building styles, textures, and vegetation, Town13 is not included in the train set to evaluate the generalization performance of trained perception models and expose overfitting issues.
Map Train Validation Test
Town01 8 2 2
Town02 8 2 2
Town03 20 4 4
Town04 20 4 4
Town05 20 4 4
Town06 20 4 4
Town07 20 4 4
Town10HD 20 4 4
Town12 48 8 8
Town13 0 8 8
Town15 36 6 6
Total 220 50 50
Table 2: Distribution of SimBEV dataset scenes across all available CARLA maps.
To create the SimBEV dataset, we used every sensor supported by SimBEV (RGB, semantic segmentation, instance segmentation, depth, and optical flow cameras; regular and semantic lidar; radar; GNSS; and IMU), positioned on the ego vehicle (a 2016 Ford Mustang) as shown in Fig. 4. We used a time step of 50 ms (i.e. 20 FPS), with each scene lasting 16 s (320 frames). In total, the SimBEV dataset contains 102,400 annotated frames, 8,315,935 3D bounding boxes (3,792,499 of which are valid), and 2,793,491,357 BEV ground truth labels. With 81.2 3D bounding boxes per frame (37.0 valid boxes per frame) it is on par with, if not surpassing, existing driving datasets [2, 4, 63, 62, 54]. A collection of front camera images displayed in Fig. 5 highlights the diversity of the SimBEV dataset.
Figure 4: Position and FoV of the perception sensors used in SimBEV to create the SimBEV dataset.
Figure 5: Front camera image samples from the SimBEV dataset.
More information about the SimBEV dataset, including sensor properties, SimBEV parameters, and dataset statistics can be found in the Supplementary Material.
5 Evaluation and Analysis
The SimBEV dataset supports a variety of tasks including 2D and 3D segmentation, depth and optical flow estimation, tracking, prediction, and localization. Here, however, we focus on the two most common tasks: BEV segmentation and 3D object detection.
5.1 Tasks and metrics
We use the intersection over union (IoU) metric for the BEV segmentation task, where for each class, a prediction is considered positive if its probability (score) is above a certain threshold. Although we calculate the IoU for each class for thresholds from 0.1 to 0.9 in increments of 0.1, we only report the values for a 0.5 threshold here. Complete results can be found in the Supplementary Material.
Our 3D object detection task metrics are inspired by [2], and rely on the Average Precision (AP) metric. We use two different methods to define a match between a predicted 3D bounding box and a ground truth one. In the first, two boxes are considered a match when their 3D IoU is above a certain threshold [12, 6]. In the second, two boxes are considered a match if the distance between their centers is below a certain threshold. As explained in [2], small translation errors for small objects (such as pedestrians) result in a low or even zero IoU value, which makes comparing the performance of camera-only models that tend to have large localization errors difficult.
For both approaches, the AP is calculated as the area under the precision-recall curve. For the first method, we use matching IoU thresholds of $\mathbb{T}=\{0.3,0.4,0.5,0.6,0.7,0.8,0.9\}$. For the second method, similar to [2], we use matching distance thresholds of $\mathbb{T}=\{0.5,1,2,4\}$ m. For both approaches we define mAP as the average over all classes and all matching thresholds:
$\mathrm{mAP}=\frac{1}{|\mathbb{T}||\mathbb{C}|}\sum_{t\in\mathbb{T},c\in\mathbb{C}}\mathrm{AP}_{t,c}.$ (1)
Similar to [2], we measure a set of True Positive metrics (TP metrics) for each predicted bounding box that is matched to a ground truth box. They are Average Translation Error (ATE), which is the Euclidean distance (in m) between box centers; Average Orientation Error (AOE), which is the smallest yaw angle difference (in rad) between the two boxes; Average Scale Error (ASE), which is equal to one minus the 3D IoU value of the two boxes after aligning for orientation and translation; and Average Velocity Error (AVE), which is the L2 norm of the difference in box velocities (in m/s). The mean TP metric (mTP) for each metric is computed by averaging over all classes and thresholds:
$\mathrm{mTP}=\frac{1}{|\mathbb{T}||\mathbb{C}|}\sum_{t\in\mathbb{T},c\in\mathbb{C}}\mathrm{TP}_{t,c}.$ (2)
Finally, similar to [2], we define the SimBEV Detection Score (SDS) as:
$\mathrm{SDS}=\frac{1}{8}\Big{(}4\,\mathrm{mAP}+\sum_{\mathrm{mTP}\in\mathbb{TP}}(1-\min(1,\mathrm{mTP}))\Big{)}.$ (3)
5.2 Evaluation results
We used BEVFusion [35], a state-of-the-art multi-sensor model for multi-task 3D perception, to evaluate both BEV segmentation and 3D object detection on the SimBEV dataset. BEVFusion offers camera-only, lidar-only, and fused (camera + lidar) model variants for each task (six variants in total), which lets us compare the performance of different sensing modalities. We trained all variants of BEVFusion using a batch size of 64, and sampled every 4th frame of the train set (17,600 samples in total). Furthermore, because of the imbalance in the number of objects for each class, we used the class-balanced grouping and sampling scheme utilized in the original implementation of BEVFusion [35, 76]. We used the validation set to tune a few hyperparameters, and evaluated the trained models on the test set. More information about BEVFusion and our training and evaluation strategy are provided in the Supplementary Material.
Modality Road Car Truck Bus Motorcycle Bicycle Rider Pedestrian Mean
C 74.1 26.7 17.0 29.5 0.0 0.0 0.0 0.0 18.4
L 89.6 75.8 77.5 84.8 44.6 4.0 38.5 27.7 55.3
C + L 89.9 77.0 79.3 84.8 45.7 4.5 37.7 32.7 56.5
Table 3: BEV segmentation IoU values (in %) for different variants of BEVFusion evaluated on the SimBEV test set.
Tab. 3 shows BEV segmentation IoU values (in %) for the three BEVFusion variants. As expected, all models achieve higher IoU values for larger objects compared to smaller ones (motorcycle, bicycle, rider, pedestrian). Moreover, IoU values for the road class (which is the only BEV segmentation class shared between SimBEV and nuScenes) are consistent with [35].
We can see from Tab. 3 that the fused model performs the best, followed closely by the lidar-only model because our lidar point cloud is very dense (approximately 225,000 points per frame). The fused model sees a 5% improvement in pedestrian IoU values, where it is helped by the additional semantic information from the camera. However, both models perform poorly when it comes to the bicycle class, though we found that the fused model achieves an IoU value of 17.9% for the bicycle class when the threshold is lowered to 0.4. It seems that because bicycles are always accompanied by a rider (and smaller and less noticeable than motorcycles), the model has difficulty distinguishing between the two and has lower confidence in its predictions.
Modality mAP mATE mAOE mASE mAVE SDS
(%) $\uparrow$ (m) $\downarrow$ (rad) $\downarrow$ $\downarrow$ (m/s) $\downarrow$ (%) $\uparrow$
C 8.0 0.313 0.561 0.100 5.35 29.3
L 36.0 0.088 0.114 0.070 4.71 52.1
C + L 36.8 0.085 0.111 0.062 4.56 52.7
Table 4: 3D object detection results for different variants of BEVFusion evaluated on the SimBEV test set using the first method.
Tab. 3 also shows that the camera-only model performs poorly (except for the road class) compared to the other two, and reports a 0 IoU value for the smaller classes. As noted before, because images lack explicit geometric information, camera-only models have difficulty localizing objects.
3D object detection results using the first and second methods are shown in Tab. 4 and Tab. 5, respectively, and a comprehensive breakdown of the results is provided in the Supplementary Material. Like before, the fused model performs the best, followed closely by the lidar-only model. We can see that the second method produces higher mAP values, which is due to its more permissive nature where, unlike the first method, two boxes can be matched even if they do not intersect at all. This permissiveness, which makes the camera-only model more comparable to the other two, can be seen when juxtaposing the mATE, mAOE, and mASE values of the two methods, with those for the second method being considerably higher. Interestingly enough, this seems to indicate that both methods result in similar SDSs. As for the large mAVE values, because none of the input data has any embedded velocity information, it is hard for the model to estimate an object’s velocity. As noted in [2], aggregating consecutive point clouds decorated by their timestamp can significantly improve mAVE.
Modality mAP mATE mAOE mASE mAVE SDS
(%) $\uparrow$ (m) $\downarrow$ (rad) $\downarrow$ $\downarrow$ (m/s) $\downarrow$ (%) $\uparrow$
C 22.7 0.708 0.686 0.131 4.92 29.8
L 46.2 0.151 0.201 0.109 4.62 54.9
C + L 46.3 0.148 0.198 0.097 4.48 55.1
Table 5: 3D object detection results for different variants of BEVFusion evaluated on the SimBEV test set using the second method.
Finally, we should note that the Swin-T [34] model used as BEVFusion’s image encoder was pretrained on ImageNet [8] and nuImages [2], which means the performance of models with image input may be less than optimal due to domain shift. Pretraining Swin-T using CARLA-based data may improve performance.
6 Conclusion
In this paper we introduce SimBEV, an extensively configurable and scalable randomized synthetic data generation tool that collects accurate BEV ground truth data and supports a variety of configurable perception and localization sensors. We hope that SimBEV enables users to rapidly create and scale different driving datasets, making it an invaluable tool for empowering researchers to explore various computer vision tasks. To showcase SimBEV, we use it to create the SimBEV dataset, a comprehensive large-scale driving dataset, which we use to benchmark BEV perception models and compare different sensing modalities and evaluation approaches. Future work will focus on expanding object classes supported by SimBEV, supporting custom sensor positions, and enabling V2X data collection.
[…]
Supplementary Material
A CARLA Simulator
SimBEV relies on a modified version of CARLA Simulator 0.9.15 [9] with an enhanced content library. Some of the improvements we made are listed below.
  • We added three new sports cars to CARLA’s vehicle library using existing 3D models [22] [^1] We used royalty-free 3D models of the three cars available on BlenderKit as the basis for the vehicles. However, the Supra and Chiron models had been removed from BlenderKit at the time of writing, so unfortunately we have no way of crediting their authors for their work.: sixth generation Ford Mustang, Toyota GR Supra, and Bugatti Chiron, shown in Fig. 6. They increase the diversity of CARLA’s vehicle library, especially when it comes to fast, high-performance vehicles. The Ford Mustang is the default vehicle for collecting SimBEV data.
  • We added lights (headlights, taillights, blinkers, etc.) to some of the older vehicle models in CARLA’s library lacking them, and redesigned existing vehicle lights in Blender using a new multi-layer approach that better visualizes modern multi-purpose lights, as shown in Fig. 7.
  • We added a set of 160 standard colors for most vehicle models to choose from (apart from a few like the firetruck), and fixed color randomization issues for a few vehicles (e.g. the bus).
  • We updated the vehicle dynamics parameters of vehicle models to better match their vehicle’s behavior and performance in the real world.
  • We added or updated pedestrian navigation information for CARLA’s Town12, Town13, and Town15 maps.
  • We updated the motorcycle and bicycle models so they select their driver models randomly each time, instead of always being assigned the same model.
  • We added lights to buildings in Town12 and fixed issues that prevented full control over building/street lights in Town12 and Town15.
SimBEV is compatible with the standard version of CARLA 0.9.15, but some features may not work properly.
Figure 6: From left to right, the Bugatti Chiron, Ford Mustang, and Toyota GR Supra vehicles added to CARLA’s vehicle library with their lights turned off (top) and on (bottom).
Figure 7: In contrast to CARLA’s segmented light design approach, our multi-layer approach can realistically visualize vehicle lights that use the same element for multiple purposes. For instance, in the Ford Mustang pictured here both position and left blinker lights are turned on.
B The SimBEV Dataset
B.1 SimBEV configuration
We configured SimBEV toward generating diverse scenarios when creating the SimBEV dataset, and collected data from every sensor supported by SimBEV (RGB, semantic segmentation, instance segmentation, depth, and optical flow cameras; regular and semantic lidar; radar; GNSS; and IMU). The configuration used for each sensor type is listed in Tab. 6 and the arrangement of sensors on the ego vehicle is detailed in Tab. 7.
Sensor type Properties
RGB camera 1600$\times$900 resolution, 80 deg FoV, $f/1.8$
All other cameras 1600$\times$900 resolution, 80 deg FoV
Lidar 128 channels, 120.0 m range, 20.0 Hz rotation frequency, 5,242,880 points per second, 10.67 to -30.67 vertical FoV,
14% general drop-off rate, 1 cm radial noise std
Semantic lidar 128 channels, 120.0 m range, 20.0 Hz rotation frequency, 5,242,880 points per second, 10.67 to -30.67 vertical FoV
Radar 120.0 m range, 100 deg horizontal FoV, 12 deg vertical FoV, 40,000 points per second
GNSS {4e-2 m, 4e-7 deg, 4e-7 deg} noise std for {altitude, latitude, longitude}
IMU 1.7e-4 rad/s gyroscope bias, {1.7e-4 m/s2, 5.6e-6 rad/s} noise std for {accelerometer, gyroscope}
Table 6: Sensor configurations used for the collection of the SimBEV dataset. std: standard deviation.
Our sensor setup was mostly inspired by [2] (e.g. the 1600$\times$900 image resolution, the arrangement of the cameras, and the lidar’s vertical FoV), though there are a few differences. Our lidars (regular and semantic) have 128 channels instead of 32 to collect a much denser point cloud, which can be downsampled by the user later on if desired. For the GNSS and IMU we used bias and noise standard deviation values of a GNSS/INS module found in a typical experimental autonomous driving platform.
Sensor $x$ (m) $y$ (m) $z$ (m) $\gamma$ (deg)
Front left camera 0.4 0.4 1.6 55
Front camera 0.6 0.0 1.6 0
Front right camera 0.4 -0.4 1.6 -55
Back left camera 0.0 0.4 1.6 110
Back camera -1.0 0.0 1.6 180
Back right camera 0.0 -0.4 1.6 -110
Left radar 0.0 1.0 0.6 90
Front radar 2.4 0.0 0.6 0
Right radar 0.0 -1.0 0.6 -90
Back radar -2.4 0.0 0.6 180
Lidar 0.0 0.0 1.8 N/A
Table 7: Arrangement of data collection sensors used in SimBEV and the SimBEV dataset. Coordinates are relative to the center of the ground plane of the ego vehicle’s 3D bounding box.
SimBEV uses the probability distributions listed in Tab. 8 to randomize the parameters that control weather in CARLA. These distributions are interdependent to ensure the resulting weather is realistic (e.g. a combination of heavy rain and clear sky is unrealistic). Each of the configured parameters is briefly discussed below.
  • Cloudiness controls the amount of clouds. Values range from 0 to 100, 0 being a clear sky.
  • Precipitation controls the intensity of rain. Values range from 0 to 100, 0 being no rain.
  • Precipitation deposits controls the amount of puddles. Values range from 0 to 100, 0 being no puddles and 100 a road filled with water.
  • Wetness controls camera image blurriness caused by rain. Values range from 0 to 100, 0 being no blurriness.
  • Wind intensity controls the strength of wind. Values range from 0 to 100, 0 being no wind at all.
  • Sun azimuth angle controls the azimuth angle of the sun. Values range from 0 to 360.
  • Sun altitude angle controls the altitude angle of the sun. Values range from -90 to 90, -90 being midnight and 90 being midday.
  • Fog density controls fog concentration or thickness. Values range from 0 to 100, 0 being no fog.
  • Fog distance controls how far away the fog starts. Values can be any nonnegative number.
  • Fog falloff controls the density of the fog (as in specific mass). Values can be any nonnegative number. If set to 0, the fog will be lighter than air and will cover the whole scene. If set to 1, the fog is approximately as dense as air. For values greater than 5 the fog will be so dense that it will be compressed to the ground level. Fog falloff is set to 0.01 in Town12, Town13, and Town15 due to their non-zero elevation.
SimBEV leaves other weather parameters (such as scattering intensity and dust storm) at their default value, though the user can change them if desired.
Parameter Symbol Distribution
Cloudiness $k_{c}$ $100\times\mathcal{B}(0.8,1.0)$
Precipitation $k_{p}$ $\mathcal{B}(0.8,0.2)\times k_{c}$ if $k_{c}>40.0$ else 0.0
Precipitation $k_{pd}$ $k_{p}+\mathcal{B}(1.2,1.6)\times(100-k_{p})$
deposits
Wetness $k_{w}$ $\min(100.0,\max(\mathcal{N}(k_{p},10.0)))$
Wind intensity $k_{wi}$ $\mathcal{U}(0.0,100.0)$
Sun azimuth $k_{az}$ $\mathcal{U}(0.0,360.0)$
angle
Sun altitude $k_{al}$ $90.0180\times\mathcal{B}(3.6,2.0)-$ 90.0
angle
Fog density $k_{f}$ $100\times\mathcal{B}(1.6,2.0)$ if $k_{c}>40.0$
or $k_{al}<10.0$ else 0.0
Fog distance $k_{fd}$ $\mathcal{LN}(3.2,0.8)$ if $k_{f}>10.0$ else 0.0
Fog falloff $k_{ff}$ $5.0\times\mathcal{B}(1.2,2.4)$ if $k_{f}>10.0$ else 1.0
Table 8: Probability distribution of SimBEV weather parameters. $\mathcal{B}$: beta distribution. $\mathcal{N}$: normal distribution. $\mathcal{U}$: uniform distribution. $\mathcal{LN}$: log-normal distribution.
Table 9 lists several other SimBEV configurations used to create the SimBEV dataset. They control various aspects of SimBEV such as scene duration, number of spawned background vehicles and pedestrians, driving behavior, chance of reckless driving, etc.
Parameter Value or distribution
Warmup duration 4 s
Scene duration 16 s
Simulation time step 50 ms
3D bounding box collection radius 120.0 m
BEV grid resolution $360\times 360$
BEV grid cell dimensions 0.4 m $\times$ 0.4 m
Distance between generated waypoints used for BEV ground truth calculation 0.4 m
Distance between generated waypoints used as spawning points 24.0 m
Number of spawned background vehicles ($s$: number of available spawn points) $\mathcal{U}_{i}(0,s-3)$
Number of spawned pedestrians $\mathcal{U}_{i}(0,640)$
Radius around the ego vehicle where background vehicles and pedestrians are spawned 400.0 m
Probability of vehicles opening door(s) when stopped 10.0%
Probability of emergency lights turned on 50.0%
Probability of ego vehicle being reckless 1.0%
Probability of other vehicles being reckless 1.0%
Minimum pedestrian speed 0.8 m/s
Maximum pedestrian speed ($r$: minimum pedestrian speed) $\max(r,\mathcal{LN}(0.16,0.64))$ m/s
Minimum street light intensity 10,000 lm
Change in street light intensity ($m$: average street light intensity of the scene) $\mathcal{U}(-m,m)$ lm
Probability of street light failure 10.0%
Maximum vehicle speed relative to the speed limit $\mathcal{U}(-20.0,40.0)\%$
Distance to front vehicle when stopped $\mathcal{N}(3.2,1.0)$ m
Traffic light green time $\mathcal{U}(4.0,28.0)$ s
Walker cross factor $\mathcal{B}(2.4,1.6)$
Table 9: SimBEV configuration used for the collection of the SimBEV dataset. $\mathcal{B}$: beta distribution. $\mathcal{N}$: normal distribution. $\mathcal{U}$: uniform distribution. $\mathcal{U}_{i}$: uniform integer distribution. $\mathcal{LN}$: log-normal distribution.
B.2 SimBEV dataset statistics
The SimBEV dataset comprises 102,400 annotated frames, 8,315,935 3D bounding boxes (3,792,499 of which are valid), and 2,793,491,357 BEV ground truth labels, broken down by class in Tab. 10. Cars and pedestrians make up the largest share of 3D bounding boxes, though they include a large number of motorcycles and bicycles as well. This makes sense given that the majority of models in CARLA’s vehicle library are cars (as opposed to e.g. only one bus model). BEV labels are dominated by the road class, followed by cars, trucks, and buses due to their larger footprint compared to the rest.
Class Total 3D Valid 3D BEV labes
bounding boxes bounding boxes
Road N/A N/A 2,674,391,899
Car 2,935,809 1,495,066 84,073,215
Truck 497,729 298,280 22,759,787
Bus 67,880 46,754 7,546,007
Motorcycle 297,132 146,083 858,136
Bicycle 214,619 100,640 187,869
Rider N/A N/A 510,521
Pedestrian 4,302,766 1,705,676 3,163,923
Total 8,315,935 3,792,499 2,793,491,357
Table 10: Breakdown of total and valid 3D bounding boxes and BEV ground truth labels by class for the SimBEV dataset.
As discussed previously, SimBEV randomizes CARLA’s weather parameters according to the distributions specified in Tab. 8. Figure 8 shows the distribution of weather across the SimBEV dataset, where precipitation (rain intensity, $k_{p}$) and fog density ($k_{f}$) values for each scene are categorized into none ($<$10%), low (10 - 40%), moderate (40 - 70%), and heavy (70 - 100%); while sun altitude angle ($k_{al}$) is categorized into night (-90 - 0 deg), dawn/dusk (0 - 6 deg), and day (6 - 90 deg). Figure 8 shows that SimBEV contains a good mix of different weather conditions, with either rain or fog present in around half of the scenes and nearly a quarter of the scenes occurring at night.
nonelowmoderateheavy00303030306060606090909090120120120120150150150150180180180180PrecipitationScenes
nonelowmoderateheavy0020202020404040406060606080808080100100100100120120120120140140140140Fog density
nightdawn/duskday004040404080808080120120120120160160160160200200200200240240240240Lighting
Figure 8: Distribution of weather across SimBEV dataset scenes.
[0, 40)[40, 80)[80, 120)[120, 160)[160, 200)[200, 240)[240, 280)[280, 320)[320, 360)[360, 400)[400, 440)[440, 480)[480, 520)[520, 560)[560, 600)600+001010101020202020303030304040404050505050606060607070707080808080CountScenesVehiclePedestrian
Figure 9: Distribution of the number of spawned vehicles (includes cars, trucks, buses, motorcycles, and bicycles) and pedestrians across SimBEV dataset scenes.
[0, 5)[5, 10)[10, 15)[15, 20)[20, 25)[25, 30)[30, 35)[35, 40)[40, 45)[45, 50)[50, 55)[55, 60)[60, 65)[65, 70)[70, 75)[75, 80)[80, 85)[85, 90)[90, 95)[95, 100)100+$10^{2}$$10^{3}$$10^{4}$$10^{5}$ramesCarPedestrian
[0, 5)[5, 10)[10, 15)[15, 20)$10^{0}$$10^{1}$$10^{2}$$10^{3}$$10^{4}$$10^{5}$ounding box per frameFramesTruckBusMotorcycleBicycle
Figure 10: Breakdown of the number of valid 3D object bounding boxes per frame by class across the SimBEV dataset.
[0, 10)[10, 20)[20, 30)[30, 40)[40, 50)[50, 60)[60, 70)[70, 80)[80, 90)[90, 100)100+$10^{3}$$10^{4}$$10^{5}$$10^{6}$istance (m)Number of objectsCarPedestrianTruckBusMotorcycleBicycle
Figure 11: Distribution of the distance of valid objects from the ego vehicle across the SimBEV dataset.
[0, 0.5)[0.5, 1)[1, 1.5)[1.5, 2)[2, 2.5)[2.5, 3)[3, 3.5)[3.5, 4)[4, 4.5)[4.5, 5)[5, 5.5)[5.5, 6)[6, 6.5)[6.5, 7)[7, 7.5)[7.5, 8)8+$10^{2}$$10^{3}$$10^{4}$$10^{5}$$10^{6}$umber of objectsPedestrian
[0, 3)[3, 6)[6, 9)[9, 12)[12, 15)[15, 18)[18, 21)[21, 24)[24, 27)[27, 30)30+$10^{0}$$10^{1}$$10^{2}$$10^{3}$$10^{4}$$10^{5}$$10^{6}$peed (m/s)Number of objectsCarTruckBusMotorcycleBicycle
Figure 12: Breakdown of the speed of valid objects across the SimBEV dataset.
Looking at the distribution of the number of spawned vehicles (includes cars, trucks, buses, motorcycles, bicycles) and pedestrians among SimBEV dataset scenes shown in Fig. 9, it is clear that the scenes range from relatively empty to congested and crowded. The distribution of pedestrians is supposed to be uniform, bus CARLA often spawns fewer pedestrians than requested, and the number of unspawned pedestrians grows rapidly when hundreds of pedestrians are requested. Hence there are many scenes with 240 - 320 pedestrians and very few with more than 480.
$10^{0}$$10^{1}$$10^{2}$$10^{3}$$10^{4}$$10^{5}$efer to captionNumber of pointsCarPedestrianTruckBusMotorcycleBicycle
[…]
Figure 13: Distribution of the number of lidar (top) and radar (bottom) points within valid 3D object bounding boxes with respect to distance from the ego vehicle.
(a) Road
(b) Car
(c) Truck
(d) Bus
(e) Motorcycle
(f) Bicycle
(g) Rider
(h) Pedestrian
Figure 14: Logarithmic BEV heat map of the SimBEV dataset for different classes.
Because Fig. 9 shows the total number of spawned vehicles and pedestrians, many of which may be far from the ego vehicle, it does not paint the full picture of what the ego vehicle observes. Hence in Fig. 10 we break down the number of valid 3D object bounding boxes per frame by class for the SimBEV dataset. The distribution of the bounding boxes is similar to [2], though our dataset offers a sizable number of frames with many (65+) valid car or pedestrian bounding boxes as well. As expected, due to fewer models in CARLA’s vehicle library, the vast majority of frames only include a handful of trucks, buses, motorcycles, and bicycles.
For more insight into the SimBEV dataset we turn to Fig. 11 through Fig. 13. Figure 11 shows that 3D object bounding box distances from the ego vehicle are nearly uniformly distributed for all classes, in contrast to [2], which is likely due to the higher density and range of our lidar point cloud. Figure 12 shows a reasonable speed range for all object classes, which is comparable to [2] with a few exceptions. We have a large number of running pedestrians (3+ m/s) which can act as edge cases for perception and behavior prediction algorithms. For other classes, our data is collected from both urban and highway environments (unlike [2] which only collected data from urban environments), leading to many fast-moving objects. Figure 13 shows the distribution of the number of lidar and radar points within valid 3D object bounding boxes with respect to distance. Consistent with [2], larger object bounding boxes have more points inside and the number of points for all classes decrease with increasing distance.
Finally, logarithmic BEV ground truth label heat maps of the SimBEV dataset for all classes are shown in Fig. 14. As expected, there is a concentration of road labels in the ego vehicle’s direction of travel, which also results in a concentration of labels in that region for all vehicular classes. In contrast, pedestrian labels are relatively evenly distributed.
C BEVFusion Implementation Details
All variants of BEVFusion [35] were trained on an Nvidia DGX A100 640GB node using sigmoid focal loss [29] and a batch size of 64. We used the original BEVFusion implementation’s settings to train the 3D object detection models, i.e. we employed the AdamW optimizer [38] with a weight decay factor of 0.01. The camera-only, lidar-only, and fused models used an initial learning rate of 2e-4, 1e-4, and 2e-4, respectively. A cyclic learning rate scheduler [52] was used for the first two models, and cosine annealing [37] was used for the latter. All three models were trained for 20 epochs. BEVFusion’s BEV segmentation variants were trained for 40 epochs using an initial learning rate of 3e-5 and a cyclic learning rate scheduler with a 4.0 target ratio. Input data was augmented (translated, rotated, scaled) for the training of all BEVFusion variants.
D Comprehensive Evaluation Results
Class 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
Road 64.1 69.1 71.6 73.3 74.1 73.8 72.6 70.5 66.4
Car 6.3 12.7 22.7 28.4 26.7 19.5 15.2 11.6 9.5
Truck 5.7 13.7 18.8 19.7 17.0 12.3 7.2 2.9 0.3
Bus 11.4 23.1 29.1 30.8 29.5 25.6 19.4 11.8 4.5
Motorcycle 0.4 2.5 1.2 0.2 0.0 0.0 0.0 0.0 0.0
Bicycle 0.1 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Rider 0.3 1.0 0.2 0.1 0.0 0.0 0.0 0.0 0.0
Pedestrian 0.2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Mean 11.0 15.3 18.0 19.1 18.4 16.4 14.3 12.1 10.1
Road 52.9 58.6 68.7 87.5 89.6 89.3 87.6 84.6 79.7
Car 6.5 41.0 60.8 72.2 75.8 68.6 57.2 45.0 28.8
Truck 14.7 49.1 66.0 75.1 77.5 71.1 60.2 47.5 29.2
Bus 27.0 65.6 78.8 84.2 84.8 79.5 70.6 60.3 46.2
Motorcycle 6.0 17.4 27.9 38.6 44.6 33.9 9.1 0.3 0.0
Bicycle 2.9 7.8 13.4 12.7 4.0 0.1 0.0 0.0 0.0
Rider 5.2 14.4 24.0 36.2 38.5 13.4 0.3 0.0 0.0
Pedestrian 3.0 11.0 19.7 31.4 27.7 2.1 0.0 0.0 0.0
Mean 14.8 33.1 44.9 54.7 55.3 44.8 35.6 29.7 23.0
Road 75.9 82.9 86.7 89.0 89.9 89.4 87.8 85.2 81.1
Car 20.3 46.1 63.8 73.8 77.0 70.0 59.2 47.5 32.3
Truck 26.1 53.8 66.7 75.5 79.3 73.8 63.2 51.3 33.8
Bus 39.6 69.0 79.9 84.5 84.8 80.3 72.5 63.1 49.7
Motorcycle 5.8 18.8 29.6 39.7 45.7 34.8 9.4 0.1 0.0
Bicycle 3.3 8.7 14.6 17.9 4.5 0.1 0.0 0.0 0.0
Rider 5.4 15.2 24.6 36.2 37.7 15.4 0.9 0.0 0.0
Pedestrian 2.8 11.3 19.9 31.6 32.7 5.8 0.1 0.0 0.0
Mean 22.4 38.2 48.2 56.0 56.5 46.2 36.6 30.9 24.6
Table 11: BEV segmentation IoU values (in %) by class and IoU threshold for the camera-only variant (top), lidar-only variant (middle), and fused variant (bottom) of BEVFusion.
Class 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
Car 33.0 29.4 23.8 17.8 12.3 7.1 2.0 0.2 0.0
Truck 30.3 27.0 23.0 18.5 13.8 9.7 5.1 1.5 0.0
Bus 31.5 30.3 27.5 25.1 21.5 15.5 10.0 5.7 0.0
Motorcycle 28.5 25.7 22.0 17.0 11.7 6.4 1.0 0.1 0.0
Bicycle 23.2 19.7 16.6 11.6 6.5 2.6 0.2 0.0 0.0
Pedestrian 2.2 1.0 0.5 0.2 0.1 0.0 0.0 0.0 0.0
Mean 24.8 22.2 18.9 15.0 11.0 6.9 3.1 1.3 0.0
Car 51.4 51.2 50.7 50.3 49.7 48.7 46.1 36.1 5.1
Truck 46.0 45.9 45.7 45.6 45.3 44.7 42.6 30.9 7.3
Bus 35.6 35.4 35.4 35.2 35.0 34.4 34.3 32.1 6.7
Motorcycle 48.1 47.3 46.6 46.0 45.3 44.3 41.5 26.3 1.4
Bicycle 56.5 56.5 56.2 55.8 55.1 52.2 42.3 23.9 0.9
Pedestrian 57.4 56.5 54.8 51.9 46.5 36.7 19.6 3.9 0.1
Mean 49.2 48.8 48.2 47.4 46.1 43.5 37.7 25.5 3.6
Car 51.5 51.3 50.8 50.4 49.8 48.9 46.5 37.2 7.2
Truck 46.2 46.0 45.8 45.7 45.3 44.7 43.4 34.0 11.4
Bus 35.6 35.6 35.5 35.4 35.2 34.9 34.3 32.1 11.8
Motorcycle 48.5 47.9 47.2 46.5 45.8 44.9 42.6 30.3 2.3
Bicycle 56.4 56.3 56.1 55.7 54.9 52.3 43.4 27.1 1.6
Pedestrian 57.3 56.6 54.7 52.0 46.7 37.7 20.8 4.5 0.1
Mean 49.2 48.9 48.4 47.6 46.3 43.9 38.5 27.5 5.7
Table 12: 3D object detection AP values (in %) by class and IoU threshold for the first method (3D IoU-based) for the camera-only variant (top), lidar-only variant (middle), and fused variant (bottom) of BEVFusion.
Class 0.5 1.0 2.0 4.0
Car 10.5 20.3 30.5 37.3
Truck 8.4 15.4 25.4 33.7
Bus 7.0 15.8 24.3 31.6
Motorcycle 16.5 26.4 31.9 34.7
Bicycle 15.6 22.1 27.7 29.7
Pedestrian 4.7 14.2 26.4 35.3
Mean 10.5 19.1 27.7 33.7
Car 34.9 45.0 51.2 51.8
Truck 21.8 36.2 45.9 46.4
Bus 34.3 34.8 35.4 35.6
Motorcycle 48.2 49.5 50.6 50.9
Bicycle 56.4 56.5 56.5 56.6
Pedestrian 38.1 52.8 59.8 60.7
Mean 38.9 45.8 49.9 50.3
Car 34.7 44.9 51.2 51.8
Truck 23.2 36.6 46.3 46.8
Bus 34.0 35.0 35.6 35.6
Motorcycle 48.0 49.1 50.2 50.8
Bicycle 56.3 56.4 56.4 56.4
Pedestrian 38.5 53.0 59.9 60.9
Mean 39.1 45.8 49.9 50.4
Table 13: 3D object detection AP values (in %) by class and distance threshold for the second method (distance-based) for the camera-only variant (top), lidar-only variant (middle), and fused variant (bottom) of BEVFusion.
BEV segmentation IoU values (in %) by class and IoU threshold for the three BEVFusion variants are shown in Tab. 11, and 3D object detection AP values (in %) by class and threshold are shown in Tab. 12 and Tab. 13 for the first and second methods, respectively. As discussed before, the biggest takeaway from the results is that the camera-only variants (for both BEV segmentation and 3D object detection) perform worse than the lidar-only and fused variants. Another takeaway is that compared to the first method, the second method for evaluating 3D object detection results allows for better comparison of the camera-only model results to that of the other two models, especially when it comes to smaller objects.